June 6, 1999 Owl shift summary

Valeri Jejer (jejer@fnal.gov)
Mon, 21 Jun 1999 07:51:25 -0500 (CDT)

Shift Summary for owl shift, 21/06/99

On Shift: Valeri, Hogan

Hours of beam delivered at all intensities: 5
Hours of beam delivered at nominal intensity: 5
Hours of E832 data (special runs not included): 3.5
Hours of special runs (laser scans, muon runs...): 0.4

Accounting of lost time:

Calling and paging split experts, cleaning cache disks, rebooting
ktev2, changing tapes and doing fresh start.
========================================================================
Runs ended this shift:
Run 13672 - 1.6M events - BEAM_MU_4E12
Run 13673 - 2.5M events - BEAM_832
Run 13674 - 1.6M events - BEAM_832

Tapes brought to FCC this shift: QK(C,K,S)019-27, QKB007-9
(including tapes for online split)

========================================================================
Accesses during shift: None

========================================================================
Overview of shift:

20 min. muon run, e832 runs,

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
13672 0:00 Found evening shift crew waiting for beam with
initialized muon run and flipped NM4AN polarity
to negative.

1:05 Beam appeares.

1:20 End muon run.

13673 1:37 Start Beam_832 run.

2:45 Onsplit disks ktev1e,2e,4e are more than 95% full.
Stopped triggers and tried to get hold of various
onsplit experts and after multiple attempts JAG
replied and helped to fix the problem.

3:30 Started triggers but got DDD veto 2, after more
detailed look into this problem found that ktev2
crashed.

13674 4:10 drt_fresh_start after saving online histograms,
unloading 72% full tapes and rebooting ktev2.

4:15 Reloaded Etot maps which failed during run init.

4:40 Noticed that spill, run time and numbers of
events (all of them) are not updated in the DAQ
monitor window. Run summary is not written
as well.

13675 5:30 Beam goes away for 15 min and hoping to fix the
DAQ monitor problem we start new run. This
doesn't help.

5:50 After returning for a very few spill beam goes
away
again due to tevatron F1 quadrupole corrector
troubles, this time for 20 min (estimate).

6:00 Online histograms show that HA HV is low, 8
channels
are below the lower limit.

6:30 Beam is back and we are running into day shift.

========================================================================
Follow up Notes on Previous Shift

* 2V TDC shows a long low-side tail, not seen in '97 running,
because the scale and number of bins there increased so that
gaussian fit doesn't work as it used to.

========================================================================
Items to Follow up on Next Shift

1) DAQ Monitoring doesn't update run, spill, time and # of events
information, run summary is not written for the current and previous
runs and these runs won't appear in ktstatus as well.

2) SDAQ Epicure button is red showing that NM2EU is off, but it's OK
according to the Epicure window.